OpenStudy (anonymous):

6/13 divided by 9/20

OpenStudy (bloomlocke367):

do you know how to divide fractions? use the "keep it, change it, flip it" method. do you know what I'm taking about?

OpenStudy (anonymous):

so 6/13*20/9

OpenStudy (anonymous):

timed exam!!!

OpenStudy (bloomlocke367):

yes!

OpenStudy (bloomlocke367):

\[\frac{ 6 }{ 13 }\times \frac{ 20 }{ 9 }=?\]

OpenStudy (bloomlocke367):

just multiply straight across

OpenStudy (anonymous):

6*20=? ------- 13*9=?

OpenStudy (jhannybean):

Remember, \[\huge \dfrac{\frac{a}{b}}{\frac{c}{d}} = \dfrac{a}{b}\cdot \frac{d}{c}\]